# Convergent Privacy Framework for Multi-layer GNNs through Contractive Message Passing

### v2

@@ A-I3 Privacy analysis for iterations
Prior works analyze cumulative privacy costs by composing iteration-wise privacy guarantees using calibrated noise and privacy composition theorems. These theorems enable modular analysis of complex algorithms by controlling the total privacy budget. Feldman et al. [focs/FeldmanMTT18, stoc/FeldmanKT20] address privacy analysis for gradient computations over a single training epoch under smooth and convex loss functions. Recent advancements [nips/ChourasiaYS21, nips/AltschulerT22, siamcomp/AltschulerBT24] improve privacy analysis for convex and strongly convex losses by eliminating dependence on infinite iterations, leveraging the Rényi DP framework. While effective, Rényi DP is lossy. $f$-DP framework offers tighter analysis through hypothesis testing curves.
For multi-hop message passing GNNs (e.g., GAP [uss/sajadmanesh2023gap]), the privacy guarantee scales linearly with the number of hops $K$. Foundational questions about privacy costs remain challenging, even for smooth convex losses over bounded domains [nips/AltschulerT22]. Recent works [nips/AltschulerT22, nips/0001S22] reveal that privacy leakage does not increase indefinitely with iterations. They demonstrate that after a brief burn-in period, additional iterations of SGD do not significantly impact privacy.
@@ A-I4 Overview of privacy attacks
Numerous privacy attacks have been developed against machine learning models, including membership inference, attribute inference, and reconstruction attacks. In the graph domain, these attacks are adapted to structured data, where the protected units are nodes, edges, or local subgraphs [uss/000100S022, sp/0011L0022, tpsisa/OlatunjiNK21]. MIA determine whether a specific node or edge appeared in the training graph, with variants studied in general ML models and specifically in GNNs. In $\mathsf{CARIBOU}$, privacy auditing module aims to empirically estimate the privacy leakage through an attacking algorithm [nips/0002NJ23, ccs/0001MMBS22]. Attribute inference attacks [olatunji2023does] aim to recover sensitive node or edge attributes from predictions. More recent graph-specific attacks further exploit message-passing behaviors to infer private structural information [uss/000100S022].
Attacking algorithms in $\mathsf{CARIBOU}$. The objective of $\mathsf{CARIBOU}$ is not to benchmark the full space of graph attacks, but to conduct a mechanism-level privacy audit of perturbed message passing under the black-box membership threat model [sp/CarliniCN0TT22, tpsisa/OlatunjiNK21] that underlies prior DP-GNN frameworks. In contrast to the upper bound obtained from theoretical privacy analysis, privacy auditing via MIA achieves the experimental true positive rate (TPR) and false positive rate (FPR) and then presents a lower bound on the privacy budget $\epsilon$ [sp/CarliniCN0TT22]. Within this setting, LinkTeller [sp/0011L0022] and the node-level GNN membership attack of Olatunji et al. [42] (G-MIA) are canonical choices: both are tailored to GNNs, operate in the transductive setting we consider, require only query access, and have open-source implementations. Using these two attacks provides a clean and reproducible evaluation of edge- and node-level leakage, enabling direct comparison with private GNNs.
Other attack families, such as attribute inference [olatunji2023does] or reconstruction attacks [zhang2022model, ijcai/ZhangLHWLLC21] require white-box knowledge. They aim to address different privacy notions or stronger threat models, and are therefore complementary rather than directly comparable to our membership-centric audit. Considering a practical and generic adversary, we do not include these attacks that require white-box access to internal embeddings or gradients, or that focus on DP notions such as attribute inference or graph-level reconstruction. Those attacks assume a strictly stronger adversary or target different sensitive objects than the edge- and node-level memberships that our DP guarantees protect. Extending our empirical evaluation to such attacks in other application scenarios is an interesting direction for future work.
@@ A-B4 Datasets
$\mathsf{CARIBOU}$ can be tested over nine datasets, including Photo and Computers [corr/abs-1811-05868], Cora and PubMed [sen2008collective], Facebook [corr/abs-1102-2166], Chain-S, Chain-M, Chain-L and Chain-X [nips/GuC0SG20].
@@ A-C1 Download the code
1git clone https://github.com/yuzhengcuhk/caribouCaemra-public.git

### v3

@@ A-I3 Privacy analysis for iterations
Prior works analyze cumulative privacy costs by composing iteration-wise privacy guarantees using calibrated noise and privacy composition theorems. These theorems enable modular analysis of complex algorithms by controlling the total privacy budget. Feldman et al. [21, 20] address privacy analysis for gradient computations over a single training epoch under smooth and convex loss functions. Recent advancements [14, 4, 3] improve privacy analysis for convex and strongly convex losses by eliminating dependence on infinite iterations, leveraging the Rényi DP framework. While effective, Rényi DP is lossy. $f$-DP framework offers tighter analysis through hypothesis testing curves.
For multi-hop message passing GNNs (e.g., GAP [59]), the privacy guarantee scales linearly with the number of hops $K$. Foundational questions about privacy costs remain challenging, even for smooth convex losses over bounded domains [4]. Recent works [4, 91] reveal that privacy leakage does not increase indefinitely with iterations. They demonstrate that after a brief burn-in period, additional iterations of SGD do not significantly impact privacy.
@@ A-I4 Overview of privacy attacks
Numerous privacy attacks have been developed against machine learning models, including membership inference, attribute inference, and reconstruction attacks. In the graph domain, these attacks are adapted to structured data, where the protected units are nodes, edges, or local subgraphs [99, 77, 53]. MIA determine whether a specific node or edge appeared in the training graph, with variants studied in general ML models and specifically in GNNs. In $\mathsf{CARIBOU}$, privacy auditing module aims to empirically estimate the privacy leakage through an attacking algorithm [64, 90]. Attribute inference attacks [51] aim to recover sensitive node or edge attributes from predictions. More recent graph-specific attacks further exploit message-passing behaviors to infer private structural information [99].
Attacking algorithms in $\mathsf{CARIBOU}$. The objective of $\mathsf{CARIBOU}$ is not to benchmark the full space of graph attacks, but to conduct a mechanism-level privacy audit of perturbed message passing under the black-box membership threat model [7, 53] that underlies prior DP-GNN frameworks. In contrast to the upper bound obtained from theoretical privacy analysis, privacy auditing via MIA achieves the experimental true positive rate (TPR) and false positive rate (FPR) and then presents a lower bound on the privacy budget $\epsilon$ [7]. Within this setting, LinkTeller [77] and the node-level GNN membership attack of Olatunji et al. [42] (G-MIA) are canonical choices: both are tailored to GNNs, operate in the transductive setting we consider, require only query access, and have open-source implementations. Using these two attacks provides a clean and reproducible evaluation of edge- and node-level leakage, enabling direct comparison with private GNNs.
Other attack families, such as attribute inference [51] or reconstruction attacks [97, 98] require white-box knowledge. They aim to address different privacy notions or stronger threat models, and are therefore complementary rather than directly comparable to our membership-centric audit. Considering a practical and generic adversary, we do not include these attacks that require white-box access to internal embeddings or gradients, or that focus on DP notions such as attribute inference or graph-level reconstruction. Those attacks assume a strictly stronger adversary or target different sensitive objects than the edge- and node-level memberships that our DP guarantees protect. Extending our empirical evaluation to such attacks in other application scenarios is an interesting direction for future work.
@@ A-B4 Datasets
$\mathsf{CARIBOU}$ can be tested over nine datasets, including Photo and Computers [62], Cora and PubMed [61], Facebook [71], Chain-S, Chain-M, Chain-L and Chain-X [25].
@@ A-C1 Download the code
1git clone https://github.com/yuzhengcuhk/caribou-public.git